OpenStudy (olivia):

SOMEONE PLEASE WHAT ARE THE X AND Y INTRCEPTS OF 3X-4Y=16

OpenStudy (olivia):

WHEN I WORKED IT I GO X (O,16/3) AND Y( 4,0)

OpenStudy (anonymous):

make x the subject 3x-4y=16 x=16/3 + (4/3)y then make y the subject

OpenStudy (anonymous):

lemmi check how its done then get back to you

OpenStudy (olivia):

SO MY ANSWER WRONG THANK YOU SO MUCH I HAVE A TEST IN AN HOUR AND NEED TO KNOW IF IM DOING IT RIGHT'

OpenStudy (anonymous):

X intercept, so its on x-axis, so y=0, then x=16/3 ---> (16/3,0) Y intercept, so its on y-axis, so x=0, then y=-4 ---> (0,-4)

OpenStudy (olivia):

THANK YOU SO MUCH

OpenStudy (olivia):

NOW Y=1/3X-1 SOLVE IT I GOT X = 1 AND Y =-1/3

OpenStudy (anonymous):

so what's the interception points?

OpenStudy (olivia):

ALL IT SAYS IS Y=1/3X-1 THEN GRAPH IT

OpenStudy (anonymous):

you know it's a straight line equation, so you will need 2 points to draw graph. just plug any random number for x to find y, that's the 1st point. do that again to get 2nd point and connect them 2, graph done.